# Towards Identifiable Unsupervised Domain Translation: A Diversified Distribution Matching Approach

### v2

@@ 16 Conv-(C-64-3, K-1, S-1, ZP-0)
ResBlock refers to block of convolutional layers with shortcut connection and optional downsampling. Specifically, ResBlock-(C-$M$-$N$, Operation) is composed of two smaller blocks, namely, Process-(C-$M$-$N$, Operation) and Shortcut-(C-$M$-$N$, Operation). The Process-(C-$M$-$N$, Operation) block has the following layers:
@@
    IN, LeakyReLU, Conv-(C-$M$-$N$, K-3, S-1, ZP-1)
The Shortcut-(C-$M$-$N$, Operation) block consists of the following layers:
@@ F.4 Dataset Details
MNIST to Rotated MNIST (MrM). We use $60,000$ training samples of the MNIST digits (LeCun et al., 2010) that have a dimension $28\times 28$ as the ${\cal X}$-domain. For the ${\cal y}$-domain, each of the $60,000$ digits is rotated by $90$ degrees. The orders of samples are shuffled in both domains to “break” the content correspondence. Under this setting, each $\bm{x}$ has a ground-truth correspondence $\bm{y}$.
Edges to Rotated Shoes (ErS). Edges2Shoes dataset (Isola et al., 2017) consists of $49,825$ training samples. We resize the all images to have $128\times 128$ pixels. The ${\cal X}$domain corresponds to the edges of the shoes, and the ${\cal Y}$-domain corresponds to the shoes that are rotated by $90$ degrees. Like in the MrM dataset, the ground-truth correspondence is known to us, which can assist evaluation.

### v3

@@ 16 Conv-(C-64-3, K-1, S-1, ZP-0)
ResBlock refers to block of convolutional layers with shortcut connection and optional downsampling. Specifically, ResBlock-(C-$M$-NN, Operation) is composed of two smaller blocks, namely, Process-(C-$M$-NN, Operation) and Shortcut-(C-$M$-NN, Operation). The Process-(C-$M$-NN, Operation) block has the following layers:
@@
    IN, LeakyReLU, Conv-(C-$M$-NN, K-3, S-1, ZP-1)
The Shortcut-(C-$M$-NN, Operation) block consists of the following layers:
@@ F.4 Dataset Details
[…]
Edges to Rotated Shoes (ErS). Edges2Shoes dataset (Isola et al., 2017) consists of $49,825$ training samples. We resize the all images to have $128\times 128$ pixels. The $\cal X$} -domain corresponds to the edges of the shoes, and the $\cal Y$}-domain corresponds to the shoes that are rotated by $90$ degrees. Like in the MrM dataset, the ground-truth correspondence is known to us, which can assist evaluation.
@@ F.6 Additional Results
There are multiple ways to define auxiliary variables for a given UDT task. However, different choices of auxiliary information can result in different level of translation performance. For example, in the MNIST example, one can alternatively use digit shape as the alphabets of the auxiliary variable. Figure 20 shows the result of using different shapes of digits. Here we use the following attributes (with corresponding digits that has those attributes): “line” : [1,2,4,5,7], “circle” : [0,6,8,9], “curve” : [0,2,3,5,6,8,9], “vertical line” : [1, 4, 5], “horizontal line” : [2,5,4,7], “curve without loops” : [2,3,5,6,9], “only vertical line” : [1], “only horizontal line” : [2]). One can see that using digit identity as auxiliary information results in a slightly enhanced performance compared to using the digit shape as auxiliary information.
@@ Appendix G Improving Existing Methods using Diversified Distribution Matching
Figure 20: Result of using different auxiliary variable for MNIST digits to rotated MNIST digits task. Using shape attributes incur LPIPS=0.19 $\pm$ 0.08, compared to LPIPS=0.11 $\pm$ 0.08 for the digit identity as auxiliary variable.